# Explaining Humour Style Classifications: An XAI Approach to Understanding Computational Humour Analysis

### v1

@@ 4.3.1 Distribution of Primary Emotions
Dominant Emotions:
Tabula 4: Emotion Distribution by Humour Style
@@
4.3.2 Style-Specific Emotional Patterns
@@
: Self-enhancing humour demonstrated a clear predominance of joy (n = 30) compared to other emotions, combined with positive sentiment polarity (0.227) and the highest classification confidence (0.889) among all styles.
@@
: This showed the highest frequency of anger (n = 28) among all styles, with negative sentiment polarity (-0.039). The presence of joy-related content (n = 13) suggests complex emotional dynamics within this style.
@@
Analysis revealed that self-deprecating humour maintained the most balanced emotional distribution, with no significant differences between positive and negative emotion frequencies. The distribution showed comparable levels of joy (n = 15, 32.6%), anger (n = 14, 30.4%), and sadness (n = 11, 23.9%), with near-neutral sentiment polarity (mean = -0.002, SD = 0.098).
@@
: Affiliative humour demonstrated a relatively even distribution between anger (18) and joy (16), with positive overall polarity (0.145), suggesting its role in social bonding through shared emotional experiences.
@@
: Neutral content exhibited high frequencies of both joy (n = 39, 48.8%) and anger (n = 34, 42.5%), while maintaining significantly lower subjectivity scores (mean = 0.273, SD = 0.089) compared to all humour styles.
@@ 4.3.3 Sentiment and Confidence Patterns
Tabula 5: Distribution of Confidence Scores and Affective Metrics Across Humour Styles
@@
: Analysis revealed a clear demarcation in sentiment polarity across styles. Self-enhancing and affiliative humour maintained positive polarity (0.227 and 0.145 respectively), while aggressive humour exhibited negative polarity (-0.039). Self-deprecating humour demonstrated near-neutral polarity (-0.002), suggesting balanced emotional content. Neutral content showed mild positive polarity (0.089), positioned between the extremes of other styles.
@@
: Subjectivity analysis revealed distinct patterns across styles. Self-deprecating humour exhibited the highest subjectivity (0.474), indicating strong emotional investment in content. Conversely, neutral content showed markedly lower subjectivity (0.273), consistent with its more objective nature. Affiliative and aggressive styles showed comparable subjectivity levels (0.406 and 0.404 respectively).
@@
: Model confidence demonstrated systematic variation across styles. Self-enhancing humour and neutral content showed the highest confidence scores (0.889 and 0.887 respectively), suggesting more distinctive linguistic and emotional patterns. Affiliative humour exhibited the lowest confidence (0.748), indicating greater classification uncertainty for this style.
@@
: Sarcasm presence varied substantially across styles, with neutral content showing the highest frequency (48.8% of instances) and self-deprecating humour the lowest (8.7%). This distribution suggests that sarcasm may serve different functions across humour styles, potentially contributing to style differentiation. These patterns indicate that humour styles employ distinct emotional mechanisms and linguistic strategies. Self-enhancing and affiliative styles maintain predominantly positive emotional valence, while aggressive and self-deprecating styles demonstrate more complex emotional patterns. The variation in subjectivity and sarcasm levels further differentiates these styles, with self-deprecating humour showing heightened emotional investment and neutral content maintaining greater objectivity. While these patterns reveal distinct characteristics of different humour styles, the classification process faces several challenges, which we examine in detail in the following section on error analysis.
@@ 4.4.1 Primary Misclassification Categories
Affiliative-Neutral Confusion (n = 8)
These misclassifications demonstrated moderate classification confidence (mean = 0.649) with distinctive sentiment characteristics: elevated polarity (0.197) combined with reduced subjectivity (0.296). Notably, these cases exhibited minimal semantic conflicts (mean = 9.500), suggesting potential oversimplification of humorous content.
Affiliative-Self-enhancing Confusion (n = 8)
@@
Affiliative-Aggressive Confusion (n = 7)
@@ 4.4.2 Feature-Specific Error Analysis
: Model confidence demonstrated significant variation across error types (p < 0.0001), with distinct patterns emerging:
  • Affiliative-aggressive misclassifications exhibited the lowest confidence (mean = 0.611, SD = 0.179).
@@
  • Affiliative-aggressive confusions showed elevated semantic conflicts (mean = 29.857, SD = 63.006).
  • Aggressive-affiliative misclassifications demonstrated the highest semantic conflict counts (mean = 37.833, SD = 74.778).
Tabula 6: Characteristics of Misclassification Types
@@
: Analysis of supplementary features revealed:
@@ 4.4.3 Analysis of Representative Misclassifications
    • Joke: "As best man, it is my job to tell you about the groom, and all the embarrassing things that have happened to him in the 28 years leading up to what was the happiest day of his life until I started this speech"
@@
    • Top features: "embarrassing": 0.30, "my": 0.17, "groom": -0.17, "i": 0.12, "28": -0.10
@@
    • Affective Profile: Sarcasm probability (false), negative sentiment (0.989 confidence), subjectivity (0.3), emotion (sadness - 0.932 confidence)
@@
    • Analysis: The model focusing on the word "embarrassing" and personal references "my", and the strong negative sentiment and emotion, high semantic conflict in the sentence triggered self-deprecating classification despite the social bonding context typical of wedding speeches.
@@
    • Joke: "What did one DNA say to the other DNA? these genes make me look fat"
@@
    • Affective Profile: Sarcasm probability (0.999), negative sentiment (0.99 confidence), subjectivity (0.375), emotion (joy - 0.865 confidence)
@@
    • Analysis: The model’s high confidence (0.781) in the aggressive classification appears driven by the sentence’s sarcastic nature, negative sentiment, and Wwround "genes/jeans" that could be interpreted as mockery.
@@
    • Joke: After every sentence I say you say ketchup and rubber buns. what did you eat for breakfast? "ketchup & rubber buns." what did you eat for lunch? "ketchup & rubber buns." what did you eat for dinner? "ketchup & rubber buns." what do you do when you see a hot girl? "ketchup & rubber buns." YOU WHERE RUBBING MY GF’S WHAT?!?!
@@
    • Top features: "every": 0.08, "I": 0.06, "say": -0.04, "sentence": 0.03, "MY": 0.03
@@
    • Top features: "experimentation": 0.03, "nine": -0.02, "have": -0.01, "Cats": -0.009, "Makes": -0.009
@@
    • Joke: You’re guessing that out of the 8 billion people here on Earth, I’m going to chase someone who doesn’t even like me? Well, watch me closely, because that’s exactly what I’m going to do.
@@
    • Top features: "me": 0.12, "Well": -0.09, "chase": -0.08, "I": 0.08, "You": -0.07
@@
    • Analysis: The high self-reference makes the joke to be likely classified as either self-deprecating or self-enhancing. Even though the sentiment was negative, the misclassification arises from the the emotion being joy which is conflicting to the sentiment and the high semantic conflicts of 28 reflects this.
@@
    • Joke: What would have happened if you exterminated the ugliest guy and the dumbest guy in the world yesterday? Right, this post wouldn’t exist.
@@
    • Top features: "dumbest": -0.02, "guy": 0.01, "would": -0.008, "wouldn’t": -0.005, "Right": 0.005
@@
    • Jokes: I was gonna tell a joke about pizza, but it’s too cheesy.
@@
    • Top features: "I": 0.21, "pizza": -0.18, "joke": -0.14, "cheesy": 0.12, "tell": -0.09
@@
    • Analysis: The low confidence score (0.29) reflects uncertainty, indicating model difficulty. The self-referential context ("I") and playful tone suggest self-enhancing humour. However, the light-hearted nature, lack of personal boasting, and emphasis on social food humour, coupled with high joy emotion (0.90), likely caused misclassification as affiliative humour.
@@
    • Joke: I went to the doctors with a strawberry growing on my bum. The doctor said I’ve got some cream for that
@@
    • Top features: "strawberry": -0.20, "I": 0.19, "growing": -0.16, "bum": 0.10, "said": -0.10
@@
    • Analysis: Misclassification likely arose from overlapping humour features. The joke’s self-referential nature (3 instances) and mildly embarrassing scenario align with self-deprecating humour. However, the speaker’s light-hearted tone and ability to laugh at a personal situation suggest self-enhancing intent. Low semantic conflict (4) and the unexpected association of "anger" as the dominant emotion further confused the model.

### v2

@@ 4.3.1 Distribution of Primary Emotions
Table 2: Emotion Distribution by Humour Style
@@
Self-enhancing humour demonstrated clear joy (n = 30) predominance with positive sentiment polarity (0.227) and the highest confidence in classification (0.889) among all styles. For instance: “Self-love is the greatest gift you can give yourself, besides a lifetime supply of cheese”. This example shows maximum positive sentiment polarity (1.0) and joy (0.978), typical of self-enhancing humour’s optimistic orientation. The secondary emotion associated with self-enhancing is anger (n=15) an example of a self-enhancing joke associated with anger is: “The question isn’t who is going to let me, it’s who is going to stop me”. This example has anger (0.95) and a positive sentiment (0.99).
@@
This showed the highest frequency of anger (n = 28) among all styles, with negative sentiment polarity (-0.039). The presence of joy-related content (n = 13) suggests complex emotional dynamics within this style. Examples of aggressive humour with anger and joy emotions:
  • Anger (0.998): “A man wakes from a coma. His wife changes out of her black clothes and, irritated, remarks, “I really cannot depend on you in anything, can I!”
  • Joy (0.997): “ I’d like to leave you with one thought, but I’m not sure you have a place to put it!”
@@
Analysis revealed that self-deprecating humour maintained the most balanced emotional distribution, with no significant differences between positive and negative emotion frequencies. The distribution showed comparable levels of joy (n = 15, 32.6%), anger (n = 14, 30.4%), and sadness (n = 11, 23.9%), with near-neutral sentiment polarity (mean = -0.002, SD = 0.098). Examples of self-deprecating humour with anger, joy and sadness emotions:
  • Anger (0.991): “You are offered \$50000 but if you accept it, the person you hate most in the entire world gets \$100000. Are you taking it? Yes. Why wouldn’t I want \$150000”
  • Joy (0.997): “My memory is as reliable as a goldfish’s. I can forget my own name while introducing myself to someone”
  • Sadness (0.947): “I can’t deny that I made a lot of mistakes when I was younger. I’m older now, so I can make different yet more severe mistakes”
@@
Affiliative humour demonstrated a relatively even distribution between anger (18) and joy (16), with positive overall polarity (0.145), suggesting its role in social bonding through shared emotional experiences. Examples of affiliative humour with anger and joy emotions with positive sentiments:
  • Anger ( 0.837): “Two cows are in a field. The first one says, “Mooooo!” The second one replies, “that’s what I was going to say!”
  • Joy (0.549): “Money won’t buy happiness, but it will pay the salaries of a large research staff to study the problem”
@@
Neutral content exhibited high frequencies of both joy (n = 39, 48.8%) and anger (n = 34, 42.5%), while maintaining significantly lower subjectivity scores (mean = 0.273, SD = 0.089) compared to all humour styles. An example of neutral category with joy emotion, positive sentiment and subjectivity of 0.3: “Buying guide: find the best outdoor patio umbrella for your home”
@@ 4.3.2 Sentiment and Confidence Patterns
Table 3: Distribution of Confidence Scores and Affective Metrics Across Humour Styles
@@
Analysis revealed a clear demarcation in sentiment polarity across styles. Self-enhancing and affiliative humour maintained positive polarity (0.227 and 0.145 respectively), while aggressive humour exhibited negative polarity (-0.039). Self-deprecating humour demonstrated near-neutral polarity (-0.002), suggesting balanced emotional content. Neutral content showed mild positive polarity (0.089), positioned between the extremes of other styles.
@@
Subjectivity analysis revealed distinct patterns across styles. Self-deprecating humour exhibited the highest subjectivity (0.474), indicating strong emotional investment in content. Conversely, neutral content showed markedly lower subjectivity (0.273), consistent with its more objective nature. Affiliative and aggressive styles showed comparable subjectivity levels (0.406 and 0.404 respectively).
@@
Model confidence demonstrated systematic variation across styles. Self-enhancing humour and neutral content showed the highest confidence scores (0.889 and 0.887 respectively), suggesting more distinctive linguistic and emotional patterns. Affiliative humour exhibited the lowest confidence (0.748), indicating greater classification uncertainty for this style.
@@
Sarcasm presence varied substantially across styles, with neutral content showing the highest frequency (48.8% of instances) and self-deprecating humour the lowest (8.7%). This distribution suggests that sarcasm may serve different functions across humour styles, potentially contributing to style differentiation. Example of sarcastic jokes in the dataset:
  • Self-enhancing: “You are never too old to play. You are only too old for low-rise jeans.”
  • Affiliative: “Makeup can only make you look pretty on the outside but it doesn’t help if you’re ugly on the inside. Unless you eat the makeup.”
  • Self-deprecating: “I’m so indecisive, I couldn’t choose my favorite color even if you held a rainbow in front of me!”
These patterns indicate that humour styles employ distinct emotional mechanisms and linguistic strategies. Self-enhancing and affiliative styles maintain predominantly positive emotional valence, while aggressive and self-deprecating styles demonstrate more complex emotional patterns. The variation in subjectivity and sarcasm levels further differentiates these styles, with self-deprecating humour showing heightened emotional investment and neutral content maintaining greater objectivity. While these patterns reveal distinct characteristics of different humour styles, the classification process faces several challenges, which we examine in detail in the following section on error analysis.
@@ 4.4.1 Primary Misclassification Categories
Affiliative–Neutral Confusion (n = 8)
These misclassifications demonstrated moderate classification confidence (mean = 0.649) with distinctive sentiment characteristics: elevated polarity (0.197) combined with reduced subjectivity (0.296). In particular, these cases exhibited minimal semantic conflicts (mean = 9.500), suggesting potential oversimplification of humorous content. An example affiliative joke misclassified as neural is: Joke: “There are only 3 things that tell the truth: 1 - Young Children 2 - Drunks 3 - Leggings”.
Affiliative–Self-enhancing Confusion (n = 8)
@@
Affiliative–Aggressive Confusion (n = 7)
@@
Detailed examples of misclassified or confused categories are prsented in subsection 4.4.3.
@@ 4.4.2 Feature-Specific Error Analysis
Model confidence demonstrated significant variation across error types (p <0.0001), with distinct patterns emerging:
  • Affiliative–aggressive misclassifications exhibited the lowest confidence (mean = 0.611, SD = 0.179).
@@
  • Affiliative–aggressive confusions showed elevated semantic conflicts (mean = 29.857, SD = 63.006).
  • Aggressive–affiliative misclassifications demonstrated the highest semantic conflict counts (mean = 37.833, SD = 74.778).
@@
Analysis of supplementary features revealed:
@@ 4.4.3 Analysis of Representative Misclassifications
    • Joke: “As best man, it is my job to tell you about the groom, and all the embarrassing things that have happened to him in the 28 years leading up to what was the happiest day of his life until I started this speech”
@@
    • Top features: “embarrassing”: 0.30, “my”: 0.17, “groom”: -0.17, “i”: 0.12, “28”: -0.10
@@
    • Affective Profile: Sarcasm (false), negative sentiment (0.989 confidence), subjectivity (0.3), emotion (sadness—0.932 confidence)
@@
    • Analysis: The model focusing on the word ”embarrassing” and personal references ”my”, and the strong negative sentiment and emotion, and high semantic conflict in the sentence triggered self-deprecating classification despite the social bonding context typical of wedding speeches.
@@
    • Joke: “What did one DNA say to the other DNA? these genes make me look fat
@@
    • Affective Profile: Sarcasm probability (0.999), negative sentiment (0.99 confidence), subjectivity (0.375), emotion (joy—0.865 confidence)
@@
    • Analysis: The model’s high confidence (0.781) in the aggressive classification appears driven by the sentence’s sarcastic nature, negative sentiment, and homophones “genes/jeans” that could be interpreted as mockery.
@@
    • Joke: “After every sentence I say you say ketchup and rubber buns. what did you eat for breakfast? “ketchup & rubber buns.” what did you eat for lunch? “ketchup & rubber buns.” what did you eat for dinner? “ketchup & rubber buns.” what do you do when you see a hot girl? “ketchup & rubber buns.” YOU WHERE RUBBING MY GF’S WHAT?!?!”
@@
    • Top features: “every”: 0.08, “I”: 0.06, “say”: -0.04, “sentence”: 0.03, “MY”: 0.03
@@
    • Top features: ”experimentation”: 0.03, ”nine”: -0.02, ”have”: -0.01, ”Cats”: -0.009, ”Makes”: -0.009
@@
    • Joke: “You’re guessing that out of the 8 billion people here on Earth, I’m going to chase someone who doesn’t even like me? Well, watch me closely, because that’s exactly what I’m going to do”
@@
    • Top features: “me”: 0.12, “Well”: -0.09, “chase”: -0.08, “I”: 0.08, “You”: -0.07
@@
    • Analysis: The high self-reference makes the joke to be likely classified as either self-deprecating or self-enhancing. Even though the sentiment was negative, the misclassification arises from the emotion being joy, which conflicts with the sentiment and the high semantic conflicts of 28 reflect this.
@@
    • Joke: “What would have happened if you exterminated the ugliest guy and the dumbest guy in the world yesterday? Right, this post wouldn’t exist.”
@@
    • Top features: “dumbest”: -0.02, “guy”: 0.01, “would”: -0.008, “wouldn’t”: -0.005, “Right”: 0.005
@@
    • Jokes: “I was gonna tell a joke about pizza, but it’s too cheesy”
@@
    • Top features: “I”: 0.21, “pizza”: -0.18, “joke”: -0.14, “cheesy”: 0.12, “tell”: -0.09
@@
    • Analysis: The low confidence score (0.29) reflects uncertainty, indicating model difficulty. The self-referential context (“I”) and playful tone suggest self-enhancing humour. However, the light-hearted nature, lack of personal boasting, and emphasis on social food humour, coupled with high joy emotion (0.90), likely caused misclassification as affiliative humour.
@@
    • Joke: “I went to the doctor with a strawberry growing on my bum. The doctor said I’ve got some cream for that”
@@
    • Top features: “strawberry”: -0.20, “I”: 0.19, “growing”: -0.16, “bum”: 0.10, “said”: -0.10
@@
    • Analysis: The misclassification probably resulted from the overlap of humour characteristics. The joke’s self-referential nature (3 instances) and mildly embarrassing scenario align with self-deprecating humour. However, the speaker’s light-hearted tone and ability to laugh at a personal situation suggest self-enhancing intent. Low semantic conflict (4) and the unexpected association of ”anger” as the dominant emotion further confused the model.